the sum of father and his son's age is 36 years. when the son becomes as old as his father now, the sum of their ages will be 84 years. find their present years.
Answer by MithrilKnights(8)

Make eqautions, let the age of father be f
Let the age of son be s
f+s = 36
So, f + f-s+f = 84
We get 3f - s = 84
Adding the two equations we get
120 = 4f
So, the age of the father is 30.
The age of the child is 6.